OpenStudy (anonymous):

Someone that is good with trinominals please show step by step Thank you :) Factor each trinomial below. Please show your work and check your answer. (1 point each) x2 – 8x + 15 a2 – a – 20 a2 – 5a – 20

OpenStudy (anonymous):

x^2-8x+15 = (x-3)(x-5)

OpenStudy (anonymous):

so is that the only step

OpenStudy (anonymous):

yes because you would just find the GCF of 15 that equals -8 which is -5 and -3

OpenStudy (anonymous):

oh ok Thank you so much :')

OpenStudy (anonymous):

yw a^2-a-20 =(a+4)(a-5)

OpenStudy (anonymous):

im not sure about the last one /.\

OpenStudy (anonymous):

its ok, thank you so much god bless you :p

OpenStudy (anonymous):

lol your welcome